OpenStudy (anonymous):

Why does expression (-2x - 6)2 can be rewritten as [-2(x + 3)]2, which equals 4(x + 3]2. Please explain where all these numbers come from.

OpenStudy (chaise):

The expression: 2(-2x-6) can be rewritten as 2(-2(x+3). This is done by factoring out a common term of two. 2(-2x-6) = 2(-2(x+3) = -4x-12 Not sure if this is what you're asking, but, thought I might comment anyway.

OpenStudy (anonymous):

Ohhh! Thank you so much!

OpenStudy (phi):

Do you mean \( (-2x - 6)^{2} \) i.e. the expression is squared?

OpenStudy (phi):

if so, you factor out a -2 to get \( ( (-2)(x-3))^{2} = (-2)^{2} (x-3)^{2} \)